OpenStudy (anonymous):

Solve the inequality 0<2/x<3/5

OpenStudy (anonymous):

\[0<\frac{ 2 }{ x }<\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

the \(0<\frac{2}{x}\) part tells you that for sure \(x>0\) otherwise \(\frac{2}{x}\) would be negative now lets look at \(\frac{2}{x}<\frac{3}{5}\)

OpenStudy (anonymous):

solve \[\frac{2}{x}<\frac{3}{5}\] by taking the reciprocal, and changing the inequality that turns in to \[\frac{x}{2}>\frac{5}{3}\] which you solve in one step by multiplying by \(2\)

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

yw